OpenStudy (anonymous):

please help solve: log(x)+log(x-14)=log(20x)

OpenStudy (anonymous):

start with \[\log(x(x-14))=\log(20x)\] then \[x(x-14)-20x\] solve the quadratic

OpenStudy (anonymous):

no times, not plus

myininaya (myininaya):

\[\log(x)+\log(x-14)-\log(20x)=0\] \[\log(\frac{x(x-14)}{20x})=\log(1)\] => \[\frac{x(x-14)}{20x}=1 =>x^2-14x-20x=0=>x^2-34x=0=>x(x-34)=0\]

OpenStudy (anonymous):

\[x^2+14x=20x\] \[x^2-6x=0\] \[x(x-6)=0\] \[x=6\] is the only solution because you cannot take the log of 0

OpenStudy (anonymous):

@myininaya too much work!!

myininaya (myininaya):

x=6?

myininaya (myininaya):

i don't think that will work satellite

OpenStudy (anonymous):

no i put + when it should have been -

myininaya (myininaya):

oh you change the problem

myininaya (myininaya):

anyways x=34 is right

OpenStudy (anonymous):

should have been \[x^2-14x=20x\]\[x^2-34x=0\] \[x(x-34)=0\] \[x=34\]

myininaya (myininaya):

since we have x(x-34)=0 x=0 or x=34 and x=0 is not a solution x=34 is only solution

OpenStudy (anonymous):

but still no need for quotients or the log of 1

myininaya (myininaya):

satellite likes to make up stuff when he does problems

OpenStudy (anonymous):

just thought i would mention it

myininaya (myininaya):

so its not wrong to do lol

OpenStudy (anonymous):

because everyone is a critic, especially me

OpenStudy (anonymous):

thank you both of you!

myininaya (myininaya):

i think we work better together

OpenStudy (anonymous):

wonderful conversation, more dramatic than the problem itself

myininaya (myininaya):

satellite and i like it that way